### v1

@@ 5 Experiments and Discussions
In the following subsections, we will use the following designs and parameter settings. In our experimental setup, we assumed that there is at most one edge between every two nodes, and all edges share identical weights for simplicity. The spectral radius  [^2] The spectral radius of a square matrix $A$ is defined as $\rho(A)=\max_{i}|\lambda_{i}|$, where $\{\lambda_{i}\}$ are the eigenvalues of $A$. of the reservoir in our experiments is set as $\lambda_{target}=0.8$ since empirical studies [38] indicate that a spectral radius close to $0.8$ realizes a favourable trade-off between fading memory and nonlinear amplification, thereby yielding best performance for the reservoir. The setting of the spectral radius can be achieved by the following steps: firstly, for a randomly generated reservoir, we assign each edge an identity weight, for example, one, to get the initial reservoir digraph $W_{initial}$. Then we can compute the maximum eigenvalue $\lambda_{max}$ of the reservoir digraph, followed by adjusting the spectral radius of the reservoir to $\lambda_{target}$ using the following formula:
@@
In the following discussion, we perform 20 tests for each of the initialization method and dataset and take the overall average of each test index, which illustrates the general applicability of our method.
@@ 5.1 Datasets and their topology
To validate our approach and compare results from different datasets on the same scale and for the convenience of comparing topological difference, we employ the following datasets [39] with value normalized into range [0.5,0.5][-0.5,0.5]:
@@
To analyze the topological differences between datasets, we employ Time-Delay Embedding (TDE) [40] to embed the time series into higher-dimensional space, followed by computing their one-dimensional persistence diagrams (1-PDs) from constructing the alpha filtration [41] of the embedded point clouds. Time-Delay Embedding is a powerful technique used in the analysis of time series data, which transforms a single observed time series into a point cloud in high-dimensional space. This method leverages the concept of phase space reconstruction, allowing for the visualization and analysis of complex time series from limited measurements [40].
@@
The choice of the embedding dimension $d$ and the time delay $\tau$ plays a crucial role in the effectiveness of TDE. The embedding dimension determines the number of past values included in each vector, capturing the dynamics of the underlying system. Meanwhile, the time delay specifies how far apart these values are spaced in time, influencing the independence between the components of the embedded vectors. In our study, by testing a variety of parameter selections, we embed the aforementioned time series into $\mathbb{R}^{3}$ with a delay parameter of $\tau=5$, which gives clear and intuitive illustration of the topological characteristics of the embedded data. This transformation not only enhances the estimation precision of models but also reveals hidden structures and patterns in the data that are not apparent in the raw time series form. The one-dimensional persistence diagram captures the topological features of loops present in the point cloud. When a point in the 1-PD exhibits significantly greater persistence than others, this indicates both the existence of a prominent loop structure in the embedded point cloud and a strong periodic mode in the original time series. This correspondence between persistent homology features and time series periodicity provides valuable insight into the underlying dynamics of the system.
The results are illustrated in Fig. 5, where we illustrate the original time series, the embedded point clouds in $\mathbb{R}^{3}$, and the 1-d persistence diagram with confidence sets. These results reveal distinct topological characteristics: the Mackey-Glass data exhibits a significant 1-dimensional topological feature as its 1-PD has one point locating above the red area, while both the MSO and Lorenz System data display several persistent 1-dimensional features with relatively great persistence, but these points are not regarded as significant. In contrast, the NARMA data is the most obvious one that shows no significant 1-dimensional topological structures. Therefore, these four datasets have different topological features from the perspective of homology.
@@
It is widely observed that Reservoir Computing models often demonstrate superior performance when dealing with datasets exhibiting pronounced periodicity [8, 7]. This enhanced capability stems from RC’s unique architectural features, which are well-suited to capture and exploit the complex temporal dependencies inherent in periodic signals. The fixed, randomly connected “reservoir” acts as a dynamic, high-dimensional feature extractor, mapping input sequences into a rich, non-linear state space. This intrinsic non-linearity and recurrent connectivity enable the reservoir to effectively represent intricate periodic patterns and their underlying dynamics, even when such patterns are embedded within noise or non-stationary components. Furthermore, the inherent memory capacity of the reservoir allows it to retain information about past inputs over extended periods, which is crucial for learning and predicting long-term periodic cycles. According to the operation of TDE, the topology of the embedded data relates to its periodicity, hence influencing the prediction results of the reservoir. This will be shown later in the experimental results.
@@ 5.2 Initial reservoir generation and its optimization
In our experiments, we tested three types of initial construction strategies for the reservoir. The first one is constructing the reservoir randomly, which typically results in a sparse reservoir digraph matrix. Additionally, we also test the small-world reservoir constructed with the Watts–Strogatz model [42] and the scale-free reservoir produced by the Barabási–Albert (BA) model [43]. These manually designed network structures have been shown in related studies to provide superior prediction accuracy and a wider spectral radius, which provide a diverse set of benchmarks for our optimization method to validate its effectiveness.
@@
Here, we compute the orthogonality measurement (OM) defined in section.4.1 of each initialization method, comparing their digraph matrices’ OM before and after optimization. Additionally, the number of rings that are added to the digraph is also counted. This evaluation is crucial because research suggests that an increase in the orthogonality of the reservoir matrix can significantly enhance its short-term memory capacity and improve the network’s ability to distinguish and retrieve different input signals. From a topological perspective, increasing the number of rings in the reservoir digraph is equivalent to its adjacency matrix having more fundamental cycle submatrices, which is the key to improving orthogonality.
@@
In this computation, the node number of the reservoir digraph is 500, which means the digraph matrix has the size $500\times 500$. Hence, the orthogonality of each digraph matrix will not be greater than 500. In the Random configuration, we set the sparsity of the reservoir to 0.99, which means that only $1\%$ of the possible edges in the fully connected graph of all nodes, which are randomly selected, actually exist. Due to the presence of a large number of zero elements, the initial orthogonality of the matrix will not be too large. In the small-world configuration, each node was initially connected to its $k=10$ nearest neighbors on a ring lattice, after which every edge was rewired with probability $p=0.1$. In the scale-free configuration, each newly added node established $m=8$ directed links according to preferential attachment.
@@
Table 1 illustrates the comparison of the average orthogonality measurement (OM) and ring addition number from 20 repeated tests of the three different initialization methods, as well as the average optimization running time. For the Random initialization approach, the reservoir’s initial sparsity results in a significant number of zero columns. Our optimization method effectively addresses this by substantially improving the average level of orthogonality, causing the OM to approach zero. Additionally, our approach introduces numerous cycles into the reservoir digraph, strongly demonstrating that the method, based on GLMY homology theory, can indeed optimize the reservoir’s structure. Similar results were observed in the Small-world initialization cases; although the increase in cycles was less than in the Random cases, the average OM was still reduced by approximately half. In the Scale-free approach, the most cycles were added. While its initial OM was already low, our optimization still reduced it by more than half. These results, while showing variations in the original orthogonality and the number of cycles that can be added across different initialization methods, collectively validate our approach’s effectiveness in increasing the orthogonality of the matrices. We can also infer that the orthogonality enhancement will be more significant with larger reservoir matrices.
@@
Table 1: Comparison of average OM, ring addition and optimization running time across different initialization methods for $500\times 500$ reservoir digraph matrices.
@@
Original OM 0.3552 0.1866 0.0958
Optimized OM 0.1598 0.0930 0.0438
Added rings 118 76 156
@@
Figure 6: Average memory capacity (MC) tested on four datasets before and after optimization. Here we compared three initialization approaches.
Figure 7: Illustration of the prediction results on four datasets. (a)–(d) are results of the Mackey-Glass, MSO, Lorenz system, and NARMA dataset, respectively. Figures on the top are prediction results of the initial randomly generated reservoir, with black being the real data and red being the predicted values. It can be seen that after a period of time, the reservoir loses its predictive ability; Figures on the bottom are prediction results of the optimized reservoir, whose predictive ability has been significantly enhanced.
@@ 5.3 Memory capacity analysis
In the subsequent experiments, the reservoir digraph’s node number is still set as 500. For the Random initial method, the reservoir sparsity is set to 0.99 (i.e., $1\%$ connectivity), which is a typical setting in practice [16]. This configuration results in a large and sparse reservoir, which effectively leverages the greater predictive power of reservoir computing.
Now we begin introducing the results of the average memory capacity (MC) of reservoirs initialized by the three aforementioned methods, both before and after optimization, which demonstrates the improvement achieved by our approach. Here, the $k_{max}$ in Equation 3 was set to $1.4n$ (where $n$ denotes the reservoir size), following the experimental setup described in [35]. For illustrative purposes, here we test the three initialization methods on the four datasets mentioned before, and compute the average MC of each tested dataset. The results presented in Fig. 6 clearly indicate that our method enhances the MC of the reservoirs. As observed in the figure, although the initial MC values vary across different initialization methods due to their distinct reservoir architectures, our optimization approach exhibits a consistent effect on each method, leading to improvements in MC across the board. Moreover, it can be concluded that the improvement of the MC is related to the type of dataset and the initialization approach. The optimization method consistently and effectively enhances the MC across all tested datasets and initialization schemes. However, the magnitude of this improvement varies significantly. For datasets like Mackey-Glass and MSO, the optimized reservoir shows a dramatic increase in MC, suggesting that the optimization is highly effective for these specific types of time series data. In contrast, for the Lorenz and NARMA datasets, the optimization provides a more modest boost to MC. Furthermore, the choice of initialization method plays a crucial role. For example, the random and small-world initialization consistently yields well improvements of MC values for the Mackey-Glass and MSO tasks, both before and after optimization, highlighting its structural advantage in these scenarios. These results are consistent with our earlier introduction in section 4.1 that MC depends on the properties of the structure and parameters of reservoirs, and may also be influenced by the data properties.

### v2

@@ 5 Experiments and Discussions
In the following subsections, we will use the following designs and parameter settings. In our experimental setup, we assumed that there is at most one edge between every two nodes, and all edges share identical weights for simplicity. The spectral radius  [^2] The spectral radius of a square matrix $A$ is defined as $\rho(A)=\max_{i}|\lambda_{i}|$, where $\{\lambda_{i}\}$ are the eigenvalues of $A$. of the reservoir in our experiments is set as $\lambda_{target}=0.8$ since empirical studies [venayagamoorthy2009effects] indicate that a spectral radius close to $0.8$ realizes a favourable trade‑off between fading memory and nonlinear amplification, thereby yielding best performance for the reservoir. The setting of the spectral radius can be achieved by the following steps: firstly, for a randomly generated reservoir, we assign each edge an identity weight, for example, one, to get the initial reservoir digraph $W_{initial}$. Then we can compute the maximum eigenvalue $\lambda_{max}$ of the reservoir digraph, followed by adjusting the spectral radius of the reservoir to $\lambda_{target}$ using the following formula:
@@
Figure 5: Reservoir size sensitivity analysis. Here we analyze reservoir sizes ranging from 100 to 500 and calculate the orthogonality measurement (OM) before and after optimization. The results presented from left to right correspond to Random, Small-world, and Scale-free initialization, respectively.
Figure 6: Spectral analysis of a reservoir with 100 nodes, 0.99 sparsity, and a fixed spectral radius of 0.8 before and after optimization. Left: eigenvalue distribution before optimization. Middle: eigenvalue distribution after optimization. Right: singular values before (blue) and after (red) optimization.
Figure 7: Illustration of five datasets, and their Time-delay embedding results in $\mathbb{R}^{3}$ and the one-dimensional persistence diagrams. (a) Result of Mackey-Glass Data. Periodicity Index is 0.1696. (b) Result of MSO Data. Periodicity Index is 0.0183. (c) Result of Lorenz System Data. Periodicity Index is 0.0183. (d) Result of NARMA Data. Periodicity Index is 0.0269. (e) Result of ETT Data. Periodicity Index is 0.0135
5.1 Reservoir Size Sensitivity and Reservoir Dynamics Analysis
First, we examine the sensitivity of our method with respect to reservoir size. In particular, we focus on optimizing the orthogonality measurement (OM) for varying reservoir sizes (ranging from 100 nodes to 500 nodes) and initialization strategies (Random, Small-world, and Scale-free). Fig. 5 demonstrates that the proposed optimization consistently reduces OM across reservoir sizes and initialization types, with the largest absolute improvements observed for initially less-orthogonal (random and scale-free) reservoirs, while preserving the spectral-radius constraint. It is noteworthy that these results are obtained while maintaining the same spectral radius constraint ($\lambda=0.8$), thus supporting our assertion that orthogonality can be improved without sacrificing stability by adjusting the structure of rings.
Then, we illustrate how our optimization method affects the reservoir dynamics. Thanks to the size sensitivity analysis, it suffices to analyze a specific reservoir size. Here we conducted a spectral analysis on a sparse reservoir with 100 nodes, 0.99 sparsity, and a fixed spectral radius of 0.8, to isolate the structural effects. Fig. 6 shows the eigenvalue distributions and singular value decay before and after optimization while keeping the spectral radius fixed at 0.8. After optimization, the eigenvalues did not exhibit the situation where a few extremely slow modes were pushed towards the spectral radius boundary; instead, the modulus distribution became more uniform. This indicates that the optimization does not rely solely on extremely slow modes but instead enhances the overall contribution of multiple moderately decaying modes, thereby giving the state space a higher effective representational capability in multiple directions and improving memory capacity. Moreover, the post-optimization singular value spectrum is noticeably flatter in its mid-range and exhibits fewer near-zero singular values, indicating a more uniform energy distribution across state directions. Dynamically, this corresponds to a redistribution of modal contributions: more moderate-decay modes rather than a few dominant very-slow modes, which yields a richer set of approximately orthogonal state components for the linear readout to exploit.
Additionally, in our spectral analysis, the eigenvalues of the reservoir weight matrix remain strictly within the target spectral radius both before and after optimization, which indicates that the optimization does not compromise this fundamental stability and hence preserves the fading memory property. Moreover, the post-optimization eigenvalue and singular value distributions suggest a redistribution of modal contributions without inducing a few excessively slow modes. Instead, a broader set of moderate-decay modes emerges, enriching the reservoir’s state representation under stable dynamics. Such a shift enhances the reservoir’s effective memory capacity while maintaining controlled fading of older inputs, consistent with the theoretical notion of fading memory.
@@ 5.2 Datasets and their topology
To validate our approach and compare results from different datasets on the same scale and for the convenience of comparing topological difference, we employ the following datasets with value normalized into the range [0.5,0.5][-0.5,0.5]:
@@
  5. 5.
    Electricity Transformer Dataset: The Electricity Transformer Dataset (ETT) dataset is one of the widely used standard public datasets in the field of long-series time series forecasting in recent years, published and systematically organized by Zhou et al. [haoyietal-informer-2021]. This dataset records various monitoring variables such as load power and oil temperature of transformers in power systems during long-term operation. It exhibits significant seasonal and periodic components, while also containing random fluctuations, making it suitable as a forecasting benchmark for real-world industrial scenarios. We use an hourly subset of the ETT dataset, selecting 10,000 hours of monitoring data beginning on 2017/1/1. In the experiments, transformer oil temperature is used as the predicting target. We use the first 5000 steps for training and the following 5000 steps for prediction.
To analyze the periodic differences between datasets, we employ Time-Delay Embedding (TDE) [takens2006detecting] to embed the time series into higher-dimensional space, followed by computing their one-dimensional persistence diagrams (1-PDs) from constructing the alpha filtration [edelsbrunner2010computational] of the embedded point clouds. Time-Delay Embedding is a powerful technique used in the analysis of time series data, which transforms a single observed time series into a point cloud in high-dimensional space. This method leverages the concept of phase space reconstruction, allowing for the visualization and analysis of complex time series from limited measurements [takens2006detecting].
@@
The choice of the embedding dimension $d$ and the time delay $\tau$ plays a crucial role in the effectiveness of TDE. The embedding dimension determines the number of past values included in each vector, capturing the dynamics of the underlying system. Meanwhile, the time delay specifies how far apart these values are spaced in time, influencing the independence between the components of the embedded vectors. In our study, by testing a variety of parameter selections, we embed the aforementioned time series into $\mathbb{R}^{3}$ with a delay parameter of $\tau=5$, which gives a clear and intuitive illustration of the topological characteristics of the embedded data. This transformation not only enhances the estimation precision of models but also reveals hidden structures and patterns in the data that are not apparent in the raw time series form. The one-dimensional persistence diagram captures the topological features of loops present in the point cloud. When a point in the 1-PD exhibits significantly greater persistence than others, this indicates both the existence of a prominent loop structure in the embedded point cloud and a strong periodic mode in the original time series [perea2015sliding]. This correspondence between persistent homology features and time series periodicity provides valuable insight into the underlying dynamics of the system. Therefore, we can use the Max Persistence in a 1-PD as a Periodicity Index (PI) to detect the periodicity of a time series:
$PI=\max{\{d_{i}-b_{i}\>|\>(b_{i},d_{i})\in PD_{1}\}}.$
In this representation, the periodic or quasi-periodic dynamics of the signal manifest as cyclic attractors [takens2006detecting, perea2015sliding]. A higher $PI$ indicates a more robust cyclic attractor in the data’s underlying dynamics, representing stronger periodicity. The results are illustrated in Fig. 7, where we illustrate the original time series, the embedded point clouds in $\mathbb{R}^{3}$, and the 1D persistence diagram with confidence sets. The title of Fig. 7 records the $PI$ of these datasets. These results reveal distinct topological characteristics: the Mackey-Glass data exhibits a significant 1-dimensional topological feature as its 1-PD has one point located above the red area, while others show no significant 1-dimensional topological structures. Hence, these five datasets differ in their prominence of periodicity.
@@
It is widely observed that Reservoir Computing models may demonstrate superior performance when dealing with datasets exhibiting pronounced periodicity [jaeger2001echo]. This enhanced capability stems from RC’s unique architectural features, which are well-suited to capture and exploit the complex temporal dependencies inherent in periodic signals. The fixed, randomly connected “reservoir” acts as a dynamic, high-dimensional feature extractor, mapping input sequences into a rich, non-linear state space. This intrinsic non-linearity and recurrent connectivity enable the reservoir to effectively represent intricate periodic patterns and their underlying dynamics, even when such patterns are embedded within noise or non-stationary components. Furthermore, the inherent memory capacity of the reservoir allows it to retain information about past inputs over extended periods, which is crucial for learning and predicting long-term periodic cycles. According to the operation of TDE, the topology of the embedded data relates to its periodicity, hence influencing the prediction results of the reservoir. This will be shown later in the experimental results.
@@ 5.3 Initial reservoir generation and its optimization
In our experiments, we tested three types of initial construction strategies for the reservoir. The first one is constructing the reservoir randomly, which typically results in a sparse reservoir digraph matrix. Additionally, we also test the small-world reservoir constructed with the Watts–Strogatz model [watts1998collective] and the scale-free reservoir produced by the Barabási–Albert (BA) model [barabasi1999emergence]. These manually designed network structures have been shown in related studies to provide superior prediction accuracy and a wider spectral radius, which provide a diverse set of benchmarks for our optimization method to validate its effectiveness.
@@
Here, we compute the orthogonality measurement (OM) defined in section 4.1 of each initialization method, comparing their digraph matrices’ OM before and after optimization. Additionally, the number of rings that are added to the digraph is also counted. This evaluation is crucial because research suggests that an increase in the orthogonality of the reservoir matrix can significantly enhance its short-term memory capacity and improve the network’s ability to distinguish and retrieve different input signals. From a topological perspective, increasing the number of rings in the reservoir digraph is equivalent to its adjacency matrix having more fundamental cycle submatrices, which is the key to improving orthogonality.
@@
In this computation, the node number of the reservoir digraph is 500, which means the digraph matrix has the size $500\times 500$. Hence, the orthogonality of each digraph matrix will not be greater than 500. In the Random configuration, we set the sparsity of the reservoir to 0.99, which means that only $1\%$ of the possible edges in the fully connected graph of all nodes, which are randomly selected, actually exist. Due to the presence of a large number of zero elements, the initial orthogonality of the matrix will not be too large. In the small-world configuration, each node was initially connected to its $k=10$ nearest neighbors on a ring lattice, after which every edge was rewired with probability $p=0.3$. In the scale-free configuration, each newly added node established $m=8$ directed links according to preferential attachment. In the following discussion, we perform 10 tests for each of the initialization methods and datasets and take the overall average of each test index, which illustrates the general applicability of our method. We compute the persistent GLMY homology and minimal representative cycles using the code provided by [dey2022efficient]. Again, we note that there is at most one edge between any two nodes, and all edges have identical weights for simplification.
@@
Table 1 illustrates the comparison of the average orthogonality measurement (OM) and ring addition number of the three different initialization methods, as well as the average optimization running time. For the Random initialization approach, the reservoir’s initial sparsity results in a significant number of zero columns. Our optimization method effectively addresses this by substantially improving the average level of orthogonality. Additionally, our approach introduces numerous cycles into the reservoir digraph, strongly demonstrating that the method, based on GLMY homology theory, can indeed optimize the reservoir’s structure. Similar results were observed in the Small-world initialization cases; although the increase in cycles was less than in the Random cases, the average OM was still reduced by approximately half. In the Scale-free approach, our optimization also reduced its OM significantly. These results, while showing variations in the original orthogonality and the number of cycles that can be added across different initialization methods, collectively validate our approach’s effectiveness in increasing the orthogonality of the matrices. We can also infer that the orthogonality enhancement will be more significant with larger reservoir matrices.
@@
Table 1: Comparison of average orthogonality measurement (OM), ring addition, and optimization running time across different initialization methods for $500\times 500$ reservoir digraph matrices.
@@
Original OM 0.3227 0.0134 0.4795
Optimized OM 0.1397 0.0074 0.2909
Added rings 112 97 88
@@
Figure 8: Illustration of the prediction results on five datasets. (a)–(e) are results of the Mackey-Glass, MSO, Lorenz system, NARMA, and ETT datasets, respectively. Figures on the top are prediction results of the initial randomly generated reservoir, with black being the real data and red being the predicted values. It can be seen that after a period of time, the reservoir loses its predictive ability; Figures on the bottom are the prediction results of the optimized reservoir, whose predictive ability has been significantly enhanced.
[…]
Mackey-Glass dataset
Random Small-world Scale-free
Origin $1.6497_{3.3122}$ $4.6585_{6.4405}$ $4.6907_{10.8505}$
GLMY Optimized $13.0021_{13.9663}$ $14.6279_{8.1599}$ $14.5722_{11.8339}$
MSO dataset
Random Small-world Scale-free
Origin $3.0975_{2.9030}$ $2.4869_{1.0932}$ $1.8996_{1.8761}$
GLMY Optimized $3.4561_{2.2308}$ $2.7297_{1.4397}$ $2.7929_{1.2467}$
Lorenz system dataset
Random Small-world Scale-free
Origin $2.7488_{2.2589}$ $2.8108_{2.3106}$ $2.0317_{0.8796}$
GLMY Optimized $5.2335_{3.8498}$ $3.7102_{2.0723}$ $3.9387_{2.5507}$
NARMA dataset
Random Small-world Scale-free
Origin $0.6518_{0.1461}$ $0.5548_{0.2050}$ $0.6869_{0.3268}$
GLMY Optimized $0.6923_{0.2479}$ $0.7996_{0.2332}$ $0.7227_{0.2506}$
ETT dataset
Random Small-world Scale-free
Origin $18.0664_{21.5056}$ $7.8563_{9.2927}$ $5.3869_{62.1683}$
GLMY Optimized $51.0430_{79.0727}$ $31.4610_{4.9896}$ $21.7803_{18.5071}$
@@ 5.4 Memory capacity analysis
In the subsequent experiments, the reservoir digraph’s node number is still set as 500. For the Random initial method, the reservoir sparsity is set to 0.99 (i.e., $1\%$ connectivity), which is a typical setting in practice [jaeger2004harnessing]. This configuration results in a large and sparse reservoir, which effectively leverages the greater predictive power of reservoir computing.
Now we begin introducing the results of the average memory capacity (MC) of reservoirs initialized by the three aforementioned methods, both before and after optimization, which demonstrates the improvement achieved by our approach. Here, the $k_{max}$ in Equation 3 was set to $1.4n$ (where $n$ denotes the reservoir size), following the experimental setup described in [farkavs2016computational]. For illustrative purposes, here we test the three initialization methods on the five datasets mentioned before, and compute the average MC and standard deviation of each tested dataset. The results presented in Table 2 clearly indicate that our method enhances the MC of the reservoirs. Although the initial MC values vary across different initialization methods due to their distinct reservoir architectures, our optimization approach exhibits a consistent effect on each method, leading to improvements in MC across the board. Moreover, it can be concluded that the improvement of the MC is related to the type of dataset and the initialization approach.
The optimization method consistently and effectively enhances the MC across all tested datasets and initialization schemes. However, the magnitude of this improvement varies significantly. For datasets like Mackey-Glass and ETT, the optimized reservoir shows a dramatic increase in MC, suggesting that the optimization is highly effective for these specific types of time series data. In contrast, for the other datasets, the optimization provides a more modest boost to MC. Furthermore, the choice of initialization method plays a crucial role. These results are consistent with our earlier introduction in section 4.1 that MC depends on the properties of the structure and parameters of reservoirs, and may also be influenced by the data properties.
@@
[…]
Mackey-Glass dataset
Random Small-world Scale-free
Origin $2.9505_{1.8449}$ $1.4882_{1.3749}$ $1.7773_{0.9896}$
Random Rewiring $3.6186_{1.3559}$ $2.3224_{1.8731}$ $3.4308_{1.4233}$
GLMY Optimized $0.3039_{0.0228}$ $0.3216_{0.0292}$ $0.3288_{0.0298}$
MSO dataset
Random Small-world Scale-free
Origin $0.6433_{0.8226}$ $0.7405_{1.0915}$ $0.7579_{0.8848}$
Random Rewiring $0.2611_{0.0855}$ $0.3496_{0.3952}$ $0.2323_{0.0119}$
GLMY Optimized $0.2248_{0.0102}$ $0.2328_{0.0164}$ $0.2216_{0.0081}$
Lorenz system dataset
Random Small-world Scale-free
Origin $1.0598_{1.2836}$ $0.9786_{1.0647}$ $0.8555_{1.0417}$
Random Rewiring $1.0706_{1.3901}$ $0.4393_{0.6146}$ $0.3919_{0.4582}$
GLMY Optimized $0.2474_{0.0075}$ $0.2454_{0.0088}$ $0.2489_{0.0088}$
NARMA dataset
Random Small-world Scale-free
Origin $0.1515_{0.0082}$ $0.1580_{0.0141}$ $0.1560_{0.0178}$
Random Rewiring $0.1539_{0.0056}$ $0.1489_{0.0028}$ $0.1530_{0.0060}$
GLMY Optimized $0.1487_{0.0028}$ $0.1475_{0.0024}$ $0.1487_{0.0035}$
ETT dataset
Random Small-world Scale-free
Origin $0.4519_{0.0665}$ $0.4438_{0.0358}$ $0.4615_{0.0362}$
Random Rewiring $0.3071_{0.1019}$ $0.1973_{0.0173}$ $0.3465_{0.0849}$
GLMY Optimized $0.1961_{0.0213}$ $0.1927_{0.0165}$ $0.2017_{0.0298}$